If the average fixed cost and average variable cost for producing 6 units of shoes are R64 and R16 respectively, what is the total cost at this level of output?
1
Expert's answer
2018-09-11T09:47:08-0400
If Q = 6 units, AFC = R64, AVC = R16, then the total cost at this level of output is:
TC = ATC*Q = (AFC + AVC)*Q = (64 + 16)*6 = R480.